# for Adaptive Streaming

### v1

@@ Abstract
The rise of Extended Reality (XR) requires efficient streaming of 3D online worlds, challenging current 3DGS representations to adapt to bandwidth-constrained environments. This paper proposes LapisGS, a layered 3DGS that supports adaptive streaming and progressive rendering. Our method constructs a layered structure for cumulative representation, incorporates dynamic opacity optimization to maintain visual fidelity, and utilizes occupancy maps to efficiently manage Gaussian splats. This proposed model offers a progressive representation supporting a continuous rendering quality adapted for bandwidth-aware streaming. Extensive experiments validate the effectiveness of our approach in balancing visual fidelity with the compactness of the model, with up to 50.71% improvement in SSIM, 286.53% improvement in LPIPS, and 318.41% reduction in model size, and shows its potential for bandwidth-adapted 3D streaming and rendering applications.
@@ 1 Introduction
Recently, 3D Gaussian Splatting (3DGS) [11] has emerged as an efficient technique for generating an explicit 3D representation from calibrated images providing photo-realistic visual quality and supporting fast rendering.
@@
While 3DGS offers advantages for representing and rendering high-fidelity 3D scenes, the resulting data size can be prohibitively large for streaming to users with various bandwidth constraints due to transfer over diverse and dynamic networks and different rendering capacities due to heterogeneous devices. Recent works [7, 5, 19, 20, 21, 14, 17] to reduce data size through model compaction or optimization are insufficient to address the dynamic nature of network conditions and device capabilities.
@@
The layered representation must meet specific requirements to effectively support adaptive 3DGS streaming. First, given subsets of Gaussian splats, the model should be able to represent complete 3D content with reduced details, allowing for the generation of a lower-quality base and enhancement layers. Second, progressive LOD should inherently share visual information with lower levels to minimize redundancy in streaming and support view-adaptive rendering. By leveraging shared visual characteristics across different levels, the progressive representation can reduce the amount of data needed to be transmitted while facilitating smooth and continuous level transition for rendering.
@@
In this paper, we propose LapisGS [^1] Lapis means ”layer” in Malay, the national language of Singapore. The logo in the title depicts kuih lapis, or ”layered cake”, a local delight in Singapore and neighboring countries, a layered representation for progressive streaming and rendering of 3DGS content. Inspired by scalable coding [24, 29, 1] and progressive LOD representation [15, 31, 36, 24, 23], this method is designed to efficiently stream and render photo-realistic 3D objects and scenes by leveraging a progressive framework that enables dynamic adaptation to varying levels of detail. As shown in Fig. 2, at the core of our approach is a layered structure for cumulative representation, where each layer adds additional details to the existing base layers, progressively refining the representation. To force coherence among layers while avoiding re-encoding lower layers, we incorporate dynamic opacity optimization during training, allowing for selective adjustment of layer contributions for optimal visual fidelity. We then utilize an occupancy map to track and exclude less important Gaussian splats during streaming and rendering, improving computational and storage efficiency.
@@
  • Extensive experiments on diverse 3D contents demonstrate the effectiveness of our method, achieving high-quality rendering and low resource cost, with up to 50.71% improvement in SSIM, 286.53% improvement in LPIPS, and 318.41% reduction in model size.
@@
To foster collaboration and further research, we will release our source code to the research community.
@@ 3.2 Continuous level transition
(a) Discrete quality adaptation.
(b) Continuous quality adaptation.
(a) Ground truth rendering.
(b) Continuous adaptive rendering.
Figure 7: The sample rendering results of Garden.
3.3 View-Adaptive Rendering
@@ 4.2 Results and Evaluation
Quantitative Comparison. We present the results from four datasets in Tabs. 4 to 4. As our comparison focuses on the trade-off between model size and rendering quality, the size and quality difference and change in size of our method compared to the Multiscale method is also presented. To better illustrate the comparison, we also show the visual quality against normalized model size for each scene and object at each resolution scale, in Fig. 8. Several key observations are highlighted.
@@
Third, LapisGS achieves notable improvements in both visual quality and model size compared to the single-scale model. At lower resolutions, our approach yields enhancements of up to 50.71% in SSIM, 286.53% in LPIPS, and a 318.41% reduction in model size. These results emphasize the necessity of constructing a multiscale 3DGS LOD model and highlight the effectiveness and efficiency of our method in maintaining visual fidelity while reducing data overhead.
@@
Figure 9: Sample renderings of Garden from Mip-NeRF360 dataset [2] at different scales.
Qualitative Comparison. We present two examples of the qualitative comparison with the other methods in Fig. 9 and Fig. 10. More detailed qualitative comparisons are shown in the supplementary materials due to the page limit.
As observed, LapisGS successfully avoids the common artifacts seen in downsampled and single-scale models, achieving results comparable to the Multiscale method but with significantly reduced model size. This improvement highlights the effectiveness of our layered progressive model in capturing fine-grained features while maintaining a compact model size, making it an ideal choice for adaptive 3DGS streaming.
@@
We present the average results in Tab. 5 and the sample renderings in Fig. 11. The per-scene quantitative and qualitative results are shown in the supplementary materials. The results highlight that without opacity optimization, model size increases significantly, and visual quality deteriorates. The visual quality degrades because lower-layer splats, which cannot capture high-frequency features, continue to influence higher-layer representations. Consequently, the model size increases significantly as additional splats are added to compensate for the deficiencies of these lower-layer splats, leading to redundancy. In contrast, our method achieves more compact models and superior visual fidelity by efficiently pruning less significant splats.
@@ 5 Discussion and Conclusion
In this paper, we introduced LapisGS, a layered progressive 3DGS designed for adaptive streaming. Our approach addresses key challenges in 3D content streaming by optimizing the balance between visual quality and model size. We leverage visual information sharing across multiple LODs, which is crucial for real-time rendering and adaptive streaming where resources are limited. A key advantage of LapisGS is its progressive nature, where each layer builds upon the previous one, allowing for efficient and coherent integration of details. By selectively optimizing the opacity of splats in lower layers, we reduce redundancy and ensure that only relevant splats enhance visual fidelity, particularly for high-frequency details. This progressive framework avoids the inefficiencies of traditional models that either downsample or train separate models for each scale. Our method demonstrates superior efficiency and flexibility, achieving substantial improvements in visual quality and model compactness, as evidenced by high SSIM and low LPIPS scores across various datasets. Additionally, it supports view-adaptive streaming and rendering through dynamic pruning and interpolation of opacity based on network conditions, ensuring smooth transitions between quality levels and a seamless user experience across different devices and network environments.
While our method performs well, there are areas for improvement. First, our approach is optimized for static scenes. Extending the model to handle dynamic elements or real-time updates in changing environments is an important direction for future research, especially for more interactive applications. Second, a notable aspect not explored in this work is the evaluation of performance under fluctuating network conditions. Understanding how our LapisGS handles variable bandwidth and latency is essential. Addressing this would require simulating network conditions and viewing behaviors, designing intricate adaptive bitrate allocation algorithms, etc. These tasks are beyond the scope of this paper, which focuses on developing an adaptive streaming model. Future work could address these aspects to ensure robust performance in diverse network environments.
@@ 6 Qualitative Results
We show the qualitative results of LapisGS alongside comparison methods on a variety of scenes, including Drjohnson and Playroom in Deep Blending dataset [8], Train and Truck in Tank&Temples dataset [13], and Room in Mip-NeRF360 dataset [2], as shown in Figs. 12 to 16.

### v2

@@ Abstract
The rise of Extended Reality (XR) requires efficient streaming of 3D online worlds, challenging current 3DGS representations to adapt to bandwidth-constrained environments. This paper proposes LapisGS, a layered 3DGS that supports adaptive streaming and progressive rendering. Our method constructs a layered structure for cumulative representation, incorporates dynamic opacity optimization to maintain visual fidelity, and utilizes occupancy maps to efficiently manage Gaussian splats. This proposed model offers a progressive representation supporting a continuous rendering quality adapted for bandwidth-aware streaming. Extensive experiments validate the effectiveness of our approach in balancing visual fidelity with the compactness of the model, with up to 50.71% improvement in SSIM, 286.53% improvement in LPIPS with 23% of the original model size, and shows its potential for bandwidth-adapted 3D streaming and rendering applications. Project page: https://yuang-ian.github.io/lapisgs/
@@ 1 Introduction
Recently, 3D Gaussian Splatting (3DGS) [11] has emerged as an efficient technique for generating an explicit 3D representation from calibrated images, providing photo-realistic visual quality and supporting fast rendering.
@@
While 3DGS offers advantages for representing and rendering high-fidelity 3D scenes, the resulting data size can be prohibitively large for streaming to users with various bandwidth constraints due to transfer over diverse and dynamic networks and different rendering capacities due to heterogeneous devices. Recent works [7, 5, 19, 20, 21, 14, 17, 28] to reduce data size through model compaction or optimization are insufficient to address the dynamic nature of network conditions and device capabilities.
@@
The layered representation must meet specific requirements to effectively support adaptive streaming for 3DGS. Firstly, given subsets of Gaussian splats, the model should be able to represent complete 3D content with reduced details, allowing for the generation of a lower-quality base and enhancement layers. Secondly, progressive levels of detail should inherently share visual information with lower levels to minimize redundancy in streaming and support view-adaptive rendering. By leveraging shared visual characteristics across different levels, the progressive representation can reduce the amount of data needed to be transmitted while facilitating smooth and continuous level transition for rendering.
@@
In this paper, we propose LapisGS [^1] Lapis means ”layer” in Malay, the national language of Singapore — the host of 3DV’25. The logo in the title depicts kuih lapis, or ”layered cake”, a local delight in Singapore and neighboring countries, a layered representation for progressive streaming and rendering of 3DGS content. Inspired by scalable coding [24, 31, 1] and progressive LOD representation [15, 33, 38, 24, 23], this method is designed to efficiently stream and render photo-realistic 3D objects and scenes by leveraging a progressive framework that enables dynamic adaptation to varying levels of detail. As shown in Fig. 2, at the core of our approach is a layered structure for cumulative representation, where each layer adds additional details to the existing base layers, progressively refining the representation. To force coherence among layers while avoiding re-encoding lower layers, we incorporate dynamic opacity optimization during training, allowing for selective adjustment of layer contributions for optimal visual fidelity. We then utilize an occupancy map to track and exclude less important Gaussian splats during streaming and rendering, improving computational and storage efficiency.
@@
  • Extensive experiments on diverse 3D contents demonstrate the effectiveness of our method, achieving high-quality rendering and low resource cost, with up to 50.71% improvement in SSIM, 286.53% improvement in LPIPS with 23% of the original model size.
@@
To foster collaboration and further research, we will release our source code to the research community and make our pre-trained layered 3DGS models publicly available.
@@ 3.2 Continuous level transition
Additionally, by dynamically and continuously adjusting the influence of splats based on their opacity, LapisGS could allow for efficient view-adaptive rendering, which optimizes rendering performance by adjusting the quality levels based on the viewer’s gaze and the distance to objects. Detailed analysis and comparison, along with illustrative figures, can be found in the supplementary materials.
@@ 4.2 Results and Evaluation
Quantitative Comparison. We present the results from four datasets in Tabs. 4 to 4. As our comparison focuses on the trade-off between model size and rendering quality, the size and quality difference and change in size of our method compared to the Multiscale method is also presented. Note that we quantify the size change relative to the new model size, which expresses the reduction as a negative percentage. A more negative value indicates a greater size decrease. To better illustrate the comparison, we also show the visual quality against normalized model size for each scene and object at each resolution scale, in Fig. 6. Several key observations are highlighted.
@@
Third, LapisGS achieves notable improvements in both visual quality and model size compared to the single-scale model. At lower resolutions, our approach yields enhancements of up to 50.71% in SSIM, 286.53% in LPIPS with 23% of the original model size. These results emphasize the necessity of constructing a multiscale 3DGS LOD model and highlight the effectiveness and efficiency of our method in maintaining visual fidelity while reducing data overhead.
@@
Figure 7: Sample renderings of Garden at different levels. LapisGS eliminates the usual artifacts seen in downsampled and single-scale models, achieving comparable quality to the Multiscale method with considerably lower data size, as shown in Tab. 4.
Qualitative Comparison. We present the qualitative comparison with the other methods in Fig. 7. More detailed qualitative comparisons are shown in the supplementary materials due to the page limit. As observed, LapisGS successfully avoids the common artifacts seen in downsampled and single-scale models, achieving results comparable to the Multiscale method but with significantly reduced model size. This improvement highlights the effectiveness of our layered progressive model in capturing fine-grained features while maintaining a compact model size, making it an ideal choice for adaptive 3DGS streaming.
@@
We present the average results in Tab. 5 and per-scene quantitative and qualitative results in the supplementary. The results highlight that without opacity optimization, model size increases significantly, and visual quality deteriorates. The visual quality degrades because lower-layer splats, which cannot capture high-frequency features, continue to influence higher-layer representations. Consequently, the model size increases significantly as additional splats are added to compensate for the deficiencies of these lower-layer splats, leading to redundancy. In contrast, our method achieves more compact models and superior visual fidelity by efficiently pruning less significant splats.
@@ 5 Discussion and Conclusion
In this paper, we introduced LapisGS, a layered progressive 3DGS designed for adaptive streaming. Our approach addresses key challenges in 3D content streaming by optimizing the balance between visual quality and model size. We leverage visual information sharing across multiple LODs, which is crucial for real-time rendering and adaptive streaming where resources are limited. A key advantage of LapisGS is its progressive nature, where each layer builds upon the previous one, allowing for efficient and coherent integration of details. By selectively optimizing the opacity of splats in lower layers, we reduce redundancy and ensure that only relevant splats enhance visual fidelity, particularly for high-frequency details. This progressive framework avoids the inefficiencies of traditional models that either downsample or train separate models for each scale. Our method demonstrates superior efficiency and flexibility, achieving substantial improvements in visual quality and model compactness, as evidenced by high SSIM and low LPIPS scores across various datasets. Additionally, it supports view-adaptive streaming and rendering through dynamic pruning and interpolation of opacity based on network conditions, ensuring smooth transitions between quality levels and a seamless user experience. For a complementary analysis and comparison, including visual illustrations, we direct readers to the supplementary materials.
While our method performs well, there are areas for improvement. First, our approach is optimized for static scenes. Extending the model to handle dynamic elements or real-time updates in changing environments is an important direction for future research, especially for more interactive applications. Second, a notable aspect not explored in this work is the evaluation of performance under fluctuating network conditions. These tasks are beyond the scope of this paper, which focuses on developing an adaptive streaming model. However, we refer readers to our recent work [30], where we develop, implement, and evaluate the first DASH-based dynamic 3DGS streaming system built on our LapisGS. This system demonstrates superior performance in both live and on-demand streaming.
@@ 7 Qualitative Results
We show the qualitative results of LapisGS alongside comparison methods on a variety of scenes, including Drjohnson and Playroom in Deep Blending dataset [8], Train and Truck in Tank&Temples dataset [13], and Room and Treehill in Mip-NeRF360 dataset [2], as shown in Figs. 10 to 15.
@@
(a) Discrete quality adaptation.
(b) Continuous quality adaptation.
(a) Ground truth rendering.
(b) Continuous adaptive rendering.
Figure 9: The sample rendering results of Garden.